OpenStudy (anonymous):

Write the equation in standard form: 6 – 2(x – y) = 3x + 5 (Points : 1) 5x – 2y = 1 5x – 2y = –1 5x + 2y = 1 2x – 5y = 1

OpenStudy (nurali):

6 – 2(x – y) = 3x + 5 6-2x+2y=3x+5 6+2y-5=3x+2x 2y+1=5x 5x-2y=1

Nnesha (nnesha):

you still need help??? @kira12341234

OpenStudy (anonymous):

yea i have a couple more questions can you help?

Nnesha (nnesha):

i will try